Latex question: I’m building this wide table, and Latex keeps indenting it about two inches, so I lose the right side. How do I make it stop indenting the table? Latex question: I’m building this wide table, and Latex keeps indenting it about two inches, so I lose the right side. How do I make it stop indenting the table?

Application-Programmer Interface

Function name  Notes
PhListFormsByType() Returns a list of all forms of the given type.
PhListFormsByOwner() Returns a list of all forms assigned to this owner.
PhListFormsByValues() Returns a list of all forms of the given type and whose field/value

pairs match the field/value pairs given in the valuelist.

PhListFormTypes()  Returns a list of all types of forms currently in Phaedo.
PhNewForm()  Returns a new form of the specified type, with default values filled in.
PhCheckinForm()  Check in the designated form
PhCheckoutForm()  Check out the designated form
PhDeleteForm()  Delete the designated form from Phaedo’s records
PhGetTemplate()  Return the template for the specified type of form
PhGetForm()  Return the body of the form pointed to by the passed header.
PhRecalculate()  Given a partially-filled-out form, recalculate new values of the

form’s automatically-filled-in fields, and return a copy of the form with the new values filled in.

PhAuthenticate()  Present name/password to Phaedo for Kerberos authentication.
PhInitialize()  Perform whatever initialization is needed by Phaedo.
PhCloseup()  Perform whatever shutdown is needed by Phaedo.

Latex question: I’m building this wide table, and Latex keeps indenting it about two inches, so I lose the right side. How do I make it stop indenting the table? Latex question: I’m building this wide table, and Latex keeps indenting it about two inches, so I lose the right side. How do I make it stop indenting the table?